OpenStudy (anonymous):

If y varies jointly as x and z, and y = 20, x = 2, and z = 5, then find k, the constant of variation.PLEASE HELP!!!

OpenStudy (anonymous):

Is it 2?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

y = k*x*z y = 2*x*z ... assume k = 2 y = 2*2*5 ... plug in x = 2, z = 5 y = 20 so this confirms that k = 2

OpenStudy (anonymous):

Thank you so much:) I have another one: What is the least common denominator of the expression a/5+a/x+5? @jim_thompson5910

OpenStudy (anonymous):

Isn't it 5(x+5)?

jimthompson5910 (jim_thompson5910):

that is correct assuming you're dealing with \[\Large \frac{a}{5} + \frac{a}{x+5}\]